OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x – 5y = –3 6x – 3y = –3 Answer 2(2x + 1) – 5y = –3 2x – 5(2x + 1) = –3 2(–2x – 1) – 5y = –3 2x – 5(–2x – 1) = –3

hero (hero):

Keke, do you have any thoughts on which one it could be?

OpenStudy (anonymous):

I DONT KNOW I TRYED BREAKING IT DOWN BUT IT DIDNT WRK

OpenStudy (anonymous):

B.!

hero (hero):

Yep, Miss T has it, but she forgot to explain her steps so that Keke gets it too

OpenStudy (anonymous):

2x - y = -4 2x = y - 4 2x + 4 = y

OpenStudy (anonymous):

*5

hero (hero):

Miss T, those are your steps? I thought you would have shown it this way: In general when using substitution for systems of equations, you solve for x or y in one equation, then substitute the expression for the appropriate variable into the other equation. In this particular situation, they want you to solve for y in the second equation, and replace its equivalent expression with y in the first equation. In the second equation, solving for y yields the following: 6x – 3y = –3 3(2x-y) = -3 2x - y = -1 2x + 1= y Therefore, y equals 2x + 1 In the first equation, we replace y with 2x+1 as follows: 2x – 5y = –3 2x-5(2x+1) = -3 And that is how we arrive at the correct solution, B

OpenStudy (anonymous):

Sorry man I'm way to lazy for all that aha.

hero (hero):

You didn't really have to do all of that. You could have just posted the correct solution steps.

hero (hero):

But don't worry, I covered for you :D

OpenStudy (anonymous):

No I expected the hero to swoop in and resolve it.!

OpenStudy (anonymous):

Corney much.? Yeah I know again to lazy to come up with a decent joke.. Tisk Tisk Shake My Head.! -_-